OpenStudy (anonymous):

Calc 1 question involving mins/max. I don't have the question in front of me, so this is going from memory, but I think I have the essentials. A person is building a rectangular garden with one side being a brick wall, and the rest being fencing. The total area is 1000sqft. The brick costs $30 per ft, the fencing is $10. Find the dimensions with the least cost. This was on an exam I just took and I tried to model it four different ways and ultimately ran out of time. Please help!

OpenStudy (camper4834):

Okay so we make a formula for what it is going to cost

OpenStudy (anonymous):

Given : A = 1000sqft = xy. my attempt : y=1000/x, A = x(1000/x) I couldn't remember where to plug the 10 and 30 into.

OpenStudy (anonymous):